A projectile of mass 1.400 kg is shot straight up with an initial speed of 27.0 m/s.

(a) How high would it go if there were no air resistance?

m

(b) If the projectile rises to a maximum height of only 27.2 m, determine the magnitude of the average force due to air resistance.

   N
Answer #1

Solution)

a)

According to Conservation of Energy,

initial K.E = P.E

so, .5mv^2 = mgh

so, h = v^2/2g = 27^2/(2*9.8) = 37.19 m (Ans)

========

Part b)

Work energy principle states,

initial K.E = P.E + work done by air resistance for 27.2 m

Now,

.5*mv^2 = mg*27.2 + F*27.2

so, .5*1.4*27^2 - 1.4*9.8*27.2= F*27.2

F = 5.04 N (Ans)

==========